# Robust Stability Analysis of Positive Lur’e System with Neural Network Feedback

### v1

@@ IV Stability Radius of Lur’e Based Neural Network Feedback Loops
In this section, we introduce a sector bound for fully connected FFNNs. This formulation enables the transformation of a NN-controlled system into the framework of a Lur’e system. By doing so, we can leverage established results from Lur’e system theory to assess the robust stability of systems controlled by neural networks. Consider a Lur’e system consisting of an LTI system (1) and an FFNN in feedback loop denoted by $\pi:\mathbb{R}^{p}\mapsto\mathbb{R}^{m}$. Suppose an FFNN with $q$ layers, described by the following:
@@
$\nu^{(i)}(t)=W^{(i)}\omega^{(i-1)}(t)+b^{(i)},\quad i=1,\dots,q,$ (14b)
$\omega^{(i)}(t)=\phi^{(i)}(\nu^{(i)}(t)),\quad i=1,\dots,q,$ (14c)
$u(t)=W^{(q+1)}\omega^{(q)}(t)+b^{(q+1)},$ (14d)
where $\omega^{(i)}(t)\in\mathbb{R}^{l_{i}}$ represents the output of the $i$-th layer, with $l_{0}=p$ and $l_{q+1}=m$, dictated by the dimensions of the LTI system’s input and output. The vector $\nu^{(i)}(t)\in\mathbb{R}^{l_{i}}$ contains the pre-activation values for the $i$th layer.
@@
We now present a result regarding the sector boundedness of the FFNN $\pi(\cdot)$ as defined in (14). We assume the following properties:
@@
  • The selected activation function lies within a sector $[a_{1},a_{2}]$, where $a_{1}<a_{2}$. [^1] Note that most known NN activation functions are sector bounded in $[0,1]$ [14]. Define $c=\max(|a_{1}|,|a_{2}|)$.
@@
Consider a fully connected FFNN $\pi(\cdot)$ with $q$ layers as defined in (14) satisfying Property 1. The network accepts an input vector $z\in\mathbb{R}^{p}_{+}$ and produces an output $u(z)\in\mathbb{R}^{m}$. The given NN is sector bounded within $[\Gamma_{1},\Gamma_{2}]$, where:
@@
with $W^{(i)}$ representing the weight matrix of layer $i$ and $|W^{(i)}|_{\text{abs}}$ being the elementwise absolute value of $W^{(i)}$.
@@
Consider system (8) in which the nonlinearity $\Phi$ is realized by a fully connected FFNN $\pi:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$, satisfying the sector condition within $[\Gamma_{1},\Gamma_{2}]$, as established in (15).
Given $A+B\Gamma_{1}C$ is Metzler, and $A+B\Gamma_{2}C$ is Hurwitz, the stability radius $r_{\mathbb{f}}$ of the NN-controlled system with respect to the structured perturbation $D\Delta E$ is:
@@
This represents the maximal perturbation magnitude under which the NN-controlled system remains globally exponentially stable.
@@
Given the structure of $\Gamma_{2}$ as described in Theorem 4 and the fact that $\Gamma_{2}\geq\Gamma_{1}$, the proof follows the same steps as the proof of Theorem 3 by substituting $\Sigma_{2}$ with $\Gamma_{2}$. Since the FFNN satisfies the sector condition within $[\Gamma_{1},\Gamma_{2}]$ and $A+B\Gamma_{2}C$ is Metzler and Hurwitz, we can apply the same analysis. Finally, we can use Lemma 3 to say
@@
If the exact upper sector bound $\Gamma_{2}$ for the neural network $\pi$ is unknown or conservative, it can be refined using the exact stability radius $r_{\mathbb{f}}$ as:
$\small\|\Gamma_{2}^{\prime}\|=\frac{1}{\left\|C(A+D\Delta E)^{-1}B\right\|},$ (18)
@@
This analysis is very useful in the case of SISO systems, where .=|.$\|.\|=|.|_{\text{abs}}$, limiting the choice of $\Gamma_{2}^{\prime}$ to ±$\pm\|\Gamma_{2}^{\prime}\|$. In this methodology, one can find the destabilizing $\Delta=r_{\mathbb{f}}$, by incrementally increasing the perturbation to a NN-controlled system and record the critical perturbation. This critical amount can be replaced in (18) to calculate the tightest sector bound for the NN. This estimation provides a tighter upper bound $\Gamma_{2}^{\prime}$, potentially improving the stability margin of the NN-controlled system.
@@ V Examples
In this section, we provide illustrative examples of the stability radius analysis using the results of the previous sections. We consider two cases: (1) a Lur’e system with structured perturbations, and (2) an instance of a NN-controlled system with structured perturbations where the exact sector bound for the NN is unknown. We demonstrate that the sector bounds obtained using (15) for the NN are conservative in this case. To address this, we refine these bounds and derive the tightest sector bounds, providing a more accurate analysis.
@@ V-A Lur’e System with Structured Perturbations
Figure 2: A schematic of the static nonlinearity tailored to lie within the sector bound $[\Sigma_{1},\Sigma_{2}]$.
@@
A static nonlinear function such as $f(y)=-1.5y+0.01y^{3}+\sin(2y)$ tailored to be sector bounded in the interval $[\Sigma_{1},\Sigma_{2}]=[-2,-0.48]$ was added to the feedback loop of the system. A schematic of this static nonlinear feedback is shown in Fig. 2. The closed loop system satisfies the conditions of Theorem 3 with this $\Sigma_{1}$ and $\Sigma_{2}$. For such system, given $D=\begin{bmatrix}1&0.5&1\end{bmatrix}^{T}$, and $E=\begin{bmatrix}0.5&1&1\end{bmatrix}$, we expect the stability radius to be
@@ V-B Neural Network-Controlled System
Consider a linear system characterized by the following matrices:
@@
The system is controlled by an FFNN trained on the data of a gain feedback controller. The sector bounds for the NN controller are calculated using Theorem 4 to be $[\Gamma_{1},\Gamma_{2}]=[-0.91,0.91]$. Since $A+B\Gamma_{1}C$ is Metzler and $A+B\Gamma_{2}C$ is Hurwitz, the closed loop Lur’e system is positive and stable. Based on Theorem 5, given $D=\begin{bmatrix}1&0&0\end{bmatrix}^{T}$, and $E=\begin{bmatrix}1&0&0\end{bmatrix}$, the stability radius of the system with the NN controller is $r_{\mathbb{f}}=2.04$. Fig. 4 shows random trajectories of the system as the perturbation increases and the system loses stability.
@@
To refine $\Gamma_{2}$, we can apply the method outlined in Corollary 1 by incrementally increasing $\Delta$ until the system reaches instability. The recorded critical value of $\Delta$, at this point, can be used in Eq. (18) to compute a more accurate $\Gamma_{2}$.
This process is illustrated in Figs. 5 and 6. As shown in Fig. 5, the system becomes unstable when the perturbation reaches $\Delta=3.15$. Fig. 6 shows the output of NN for random positive inputs and the recalculated upper bound $\Gamma_{2}^{\prime}z$. It is demonstrated that the NN output now falls outside the sector bound for the calculated $\Gamma^{\prime}_{2}$ using the critical amount of $\Delta=3.15$ in Eq. (18).
Figure 5: Random output trajectories of the system for various $\Delta$ values. As illustrated here, the system becomes unstable at $\Delta=3.15$.
Figure 6: Plot of NN output vs. sector bounds for random positive inputs. Each figure shows the calculated value $\Gamma_{2}^{\prime}$ for the corresponding $\Delta$ in Fig. 5, using Eq. (18). As illustrated here, the NN output falls out of sector bound for the amount of $\Gamma_{2}^{\prime}=0.25$ which corresponds to $\Delta=3.15$.
@@ VI Discussion and Conclusion
The explicit expressions for the stability radius derived in Theorem 3 and Corollary 5 defines the robustness of positive Lur’e systems and NN-controlled systems against structured uncertainties. By leveraging positivity constraints, we utilize Aizerman’s conjecture and the stability radius of Metzler matrices to evaluate the robustness of nonlinear Lur’e systems. These results highlight several key insights:
Dependence on the Upper Sector Bound: The stability radius of nonlinear Lur’e system is determined by the worst-case (upper) sector bound ($\Sigma_{2}$ or $\Gamma_{2}$). This underscores the importance of accurately characterizing the sector in which the nonlinearity or NN operates.
@@
Applicability to Neural Networks: By modeling the NN controller as a sector-bounded nonlinearity, we bridge the gap between traditional control theory and modern NN-based controllers. The derived stability radius provides a quantitative measure of the NN-controlled system’s robustness.
Refinement of Sector Bounds: The methodology outlined in the corollary allows practitioners to iteratively improve the sector bounds for neural networks, leading to more accurate and less conservative stability assessments.

### v3

@@ IV Stability Radius of Lur’e Based Neural Network Feedback Loops
In this section, we introduce a sector bound for fully connected FFNNs. This formulation enables the transformation of an NN-controlled system into the framework of a Lur’e system. By doing so, we can leverage established results from Lur’e system theory to assess the robust stability of systems controlled by NNs. Consider a Lur’e system consisting of an LTI system (1) and an FFNN in feedback loop denoted by $\pi:\mathbb{R}^{p}\mapsto\mathbb{R}^{m}$. Suppose an FFNN with $q$ layers described by the following equations:
@@
$\omega^{(i)}(t)=\phi^{(i)}(W^{(i)}\omega^{(i-1)}(t)+b^{(i)}),\qquad i=1,\dots,q,$ (14b)
$u(t)=W^{(q+1)}\omega^{(q)}(t)+b^{(q+1)},$ (14c)
where $\omega^{(i)}(t)\in\mathbb{R}^{l_{i}}$ represents the output of the $i$-th layer, with $l_{0}=p$ and $l_{q+1}=m$, dictated by the dimensions of the LTI system’s input and output.
@@
We now present a result regarding the sector bounds of general FFNNs as in (14). We assume the following properties:
@@
  • The selected activation function lies within a sector $a_{1},a_{2}]$, where $a_{1}<a_{2}$ [^1] Most famous NN activation functions are sector bounded in $[0,1]$ [14].. Define $c=\max(|a_{1}|,|a_{2}|)$.
@@
Consider a fully connected FFNN $\pi(\cdot)$ with $q$ layers as defined in (14) satisfying Property 1. The network accepts an input vector $z\in\mathbb{R}^{p}_{+}$ and produces an output $u(z)\in\mathbb{R}^{m}$. The given NN is sector bounded within $\Gamma_{1},\Gamma_{2}]$, where
@@
with $|W^{(i)}|_{\text{abs}}$ being the elementwise absolute value of $W^{(i)}$.
@@
Consider system (8) in which the nonlinearity $\Phi$ is realized by a fully connected FFNN $\pi:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$, satisfying the sector condition within $\Gamma_{1},\Gamma_{2}]$, as established in (15). Given $A+B\Gamma_{1}C$ is Metzler and $A+B\Gamma_{2}C$ is Hurwitz, the stability radius $r_{\mathbb{f}}$ of the NN-controlled system with respect to the structured perturbation $D\Delta E$ is
@@
This value represents the maximal perturbation norm under which the NN-controlled system remains globally exponentially stable.
@@
Given the structure of $\Gamma_{2}$ as described in Theorem 4 and the fact that $\Gamma_{2}\geq\Gamma_{1}$, the proof follows the same steps as the proof of Theorem 3 by substituting $\Sigma_{2}$ with $\Gamma_{2}$. Since the FFNN satisfies the sector condition within $\Gamma_{1},\Gamma_{2}]$ and $A+B\Gamma_{2}C$ is Metzler and Hurwitz, we can apply the same analysis. Finally, we can use Lemma 3 to say
@@
If the exact upper sector bound $\Gamma_{2}$ for the NN $\pi$ is unknown or conservative, it can be refined using the exact stability radius $r_{\mathbb{f}}$ as
$\small\|\Gamma_{2,\text{refined}}\|=\frac{1}{\left\|C(A+D\Delta E)^{-1}B\right\|},$ (18)
@@
This analysis is very useful in the case of SISO systems, where $\|.\|=|.|_{\text{abs}}$, limiting the choice of $\Gamma_{2,\text{refined}}$ to $\pm\|\Gamma_{2,\text{refined}}\|$. In this methodology, one can find the destabilizing $\Delta=r_{\mathbb{f}}$, by incrementally increasing the perturbation to an NN-controlled system and recording the critical perturbation. This critical amount can be used in (18) to calculate the tightest sector bound for the NN. This estimation provides a tighter upper bound $\Gamma_{2,\text{refined}}$, potentially improving the stability margin of the NN-controlled system.
@@ V Examples
In this section, we provide illustrative examples of the stability radius analysis using the results of the previous sections. We consider two cases: (1) a Lur’e system with structured uncertainty and (2) an instance of an NN-controlled system with structured uncertainty where the exact sector bound for the NN is unknown. We demonstrate that the sector bounds obtained using (15) for the NN are conservative in this case. To address this, we refine these bounds and derive the tightest sector bounds, providing a more accurate analysis.
@@ V-A Lur’e System with Structured Perturbations
A static nonlinear function $f(y)=-1.5y+0.01y^{3}+\sin(2y)$ tailored to be sector bounded in the interval $\Sigma_{1},\Sigma_{2}]=[-2,-0.48]$ was added to the feedback loop of the system. A schematic of this static nonlinear feedback is shown in Fig. 2.
Figure 2: A schematic of the static nonlinearity tailored to lie within the sector bound $\Sigma_{1},\Sigma_{2}]$.
The closed-loop system satisfies the conditions of Theorem 3 with this $\Sigma_{1}$ and $\Sigma_{2}$. For such a system, given $D=\begin{bmatrix}1&0.5&1\end{bmatrix}^{T}$, and $E=\begin{bmatrix}0.5&1&1\end{bmatrix}$, we expect the stability radius to be
@@ V-B Neural Network-Controlled System
Consider a linear system characterized by the matrices:
@@
The system is controlled by an FFNN trained on the data of a gain feedback controller. Using Theorem 4, the NN sector bounds are calculated to be $\Gamma_{1},\Gamma_{2}]=[-0.91,0.91]$. Since $A+B\Gamma_{1}C$ is Metzler and $A+B\Gamma_{2}C$ is Hurwitz, the closed loop Lur’e system is positive and stable. Based on Theorem 5, given $D=\begin{bmatrix}1&0&0\end{bmatrix}^{T}$, and $E=\begin{bmatrix}1&0&0\end{bmatrix}$, the stability radius of the system with the NN controller is $r_{\mathbb{f}}=2.04$. Fig. 4 shows random system trajectories as the perturbation increases and the system loses stability.
@@
To refine $\Gamma_{2}$, we apply the procedure described in Corollary 1, incrementally increasing $\Delta$ until the system becomes unstable. The corresponding critical value of $\Delta$ is then substituted into equation (18) to obtain a more accurate $\Gamma_{2}$.
This process is illustrated in Figs. 5 and 6. As shown in Fig. 5, the system becomes unstable when the perturbation reaches $\Delta=3.15$. Fig. 6 shows the output of NN for random positive inputs and the recalculated upper bound $\Gamma_{2,\text{refined}}z$. It is demonstrated that the NN output now falls outside the sector bound for the calculated $\Gamma_{2,\text{refined}}$ using the critical amount of $\Delta=3.15$ in equation (18).
Figure 5: Random output trajectories of the system for various $\Delta$ values. As illustrated, the system becomes unstable at $\Delta=3.15$.
Figure 6: Plot of NN output vs. sector bounds for random positive inputs. Each figure shows the calculated value $\Gamma_{2,\text{refined}}$ for the corresponding $\Delta$ in Fig. 5, using equation (18). As illustrated here, the NN output falls out of sector bound for the amount of $\Gamma_{2,\text{refined}}=0.25$ which corresponds to $\Delta=3.15$.
@@ VI Discussion and Conclusion
By leveraging positivity constraints, we utilize the Aizerman conjecture and the stability radius of Metzler matrices to evaluate the robustness of nonlinear Lur’e systems. The explicit expressions for the stability radius derived in Theorem 3 and Theorem 5 define the robustness of positive Lur’e systems and NN-controlled systems against structured uncertainties. Corollary 1 also proposes a method for finding the tightest sector bounds for NNs. These results highlight several key insights:
Dependence on the Upper Sector Bound: The stability radius of nonlinear Lur’e systems is determined by the worst-case (upper) sector bound ($\Sigma_{2}$ or $\Gamma_{2}$). This underscores the importance of accurately characterizing the sector in which the nonlinearity or NN operates.
Limitation: Positivity is not imposed on the linear system; the matrix $A$ may be non-Metzler. Thus, the analysis covers a broad class of systems. It is only required that the closed-loop nonlinear system be positive through $A+B\Sigma_{1}C.$
NN bound refinement: Unlike other bounding methods such as CROWN and IBP, which are sound but incomplete, our sound and complete method yields an exact upper sector bound for the NN in the loop. It requires no knowledge of the NN parameters (weights, biases), treating the NN as a black box and shifting the analysis to the linear system—greatly reducing computational effort.
@@
Applicability to NNs: By modeling the NN controller as a sector-bounded nonlinearity, we bridge the gap between traditional control theory and modern NN-based controllers. The derived stability radius provides a quantitative measure for the robustness of NN-controlled systems.
Future work may include the analysis of alternative perturbation structures, the integration of additional results from positive linear systems theory into the Lur’e framework, and incorporation of other NN architectures in the feedback loop.